The return on the Rush Corporation in the state of recession is estimated to be-24% and the return on Rush in the state of boom is estimated to be 35%. The return on the Oberman Corporation in the state of recession is estimated to be 45% and the return on Oberman in the state of boom is estimated to be-17%. Given this information, what is the covariance between Rush and Oberman if there is a 0.50 probability that the economy will be in the state of boom and a 0.50 probability that the economy will be in the state of recession Place your answer in decimal form and not as a percentage.
Answer #1

Expected Return for Rush Corporation = 0.5*-24% + 0.5* 35% = 4.5%
Expected Return for Oberman Corporation = 0.5*45%+0.5*-17% = 14%

Covariance formula = Probability of Boom *(Return of Rush Corporation in boom - Expected return of Rush)*(Return of Oberman Corporation in boom - Expected return of Oberman )+ Probability of Recession*(Return of Rush Corporation in boom - Expected return of Rush)*(Return of Oberman Corporation in boom - Expected return of Oberman ) = 0.5*(35%-4.5%)*(-17%-14%)+0.5*(-24%-4.5%)*(45%-17%) = -0.0872

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well